OpenStudy (rane):

Take the reaction: NH3 + O2 NO + H2O. In an experiment, 3.25 g of NH3 are allowed to react with 3.50 g of O2. a. Which reactant is the limiting reagent? b. How many grams of NO are formed?

OpenStudy (rane):

can someone pls help. i have an exam tomorrow

OpenStudy (rane):

@.Sam. @AravindG @ganeshie8

sam (.sam.):

Convert grams of NH3 and grams of O2 into moles, then see which is larger

OpenStudy (rane):

i have found my limited reagent bt just confused abt 2nd part where i have to find mass

sam (.sam.):

Well since the mole ratio of O2 and NO is 1, so n(O2)=n(NO)=0.1094mol

sam (.sam.):

Then just multiply 0.1094mol by its molar mass and you'll get the mass

sam (.sam.):

its molar mass=NO's molar mass

OpenStudy (rane):

so.. is the and 6.0928

OpenStudy (rane):

ans*

sam (.sam.):

Molar mass of NO=14+16=30g/mol Using 0.1094mol, \[0.1094mol \times 30g/mol=3.282g~of~NO\]

OpenStudy (rane):

sory is it 5.7

OpenStudy (rane):

but i get 5.7 if i do this

sam (.sam.):

Did you use a calculator, lol

OpenStudy (rane):

um..yess lol

sam (.sam.):

Yeah just multiply 0.1094 by 30

OpenStudy (rane):

i did

OpenStudy (rane):

anyway thank u for showing me how to do it

sam (.sam.):

yw :)